What is the shortest distance from the surface xy+12x+z2=144 to the origin?

distance=?
1
Expert's answer
2020-04-07T12:08:32-0400

Let the square of the distance function be "f(x,y,z)=x^2+y^2+z^2"

Given, the constraint function "g(x,y,z)=xy+12x+z^2"


Apply Lagrange Multiplier as,


"\nabla f(x,y,z)=\lambda g(x,y,z)"


So, find the gradient both sides to obtain,


"<2x,2y,2z>=\lambda\ <y+12,x,2z>"

Equate to gradients of both sides as,


"2x=\lambda (y+12),
2y=\lambda (x),
2z=\lambda (2z)"

From equation "2z=\lambda (2z),\lambda=1" , so plug "\lambda=1" into "2y=\lambda (x)" to obtain "x=2y".

Next, plug "x=2y" and "\lambda=1" into equation "2x=\lambda (y+12)" and solve for "y" as,



"2(2y)=(y+12)""4y=y+12"

Subtract "y" from both sides to isolate variable "y" from constant as,


"4y-y=y+12-y""3y=12"

Divide both sides by "3" to obtain,


"\frac{3y}{3}=\frac{12}{3}""y=4"

So, the value of "y" is "y=4" and "x=2y=2(4)=8"


Now, plug the values of "x" and "y" into constraint equation "xy+12x+z^2=144" and solve for "z" as,


"(8)(4)+12(8)+z^2=144""32+48+z^2=144""80+z^2=144"

Subtract "80" from both sides to obtain,


"80+z^2-80=144-80""z^2=64""z=\u00b18"

Now, plug "x=8,y=4" and "z=\u00b18" into objective function "f(x,y,z)=x^2+y^2+z^2" to obtain,


"f(8,4,\u00b18)=(8)^2+(4)^2+(\u00b18)^2=64+16+64=144"

Next, find the square root of objective function as,


"d=\sqrt{144}=12"


Therefore, the shortest distance from the surface "xy+12x+z^2=144" to the origin is "[12]".